# Domain Generalizable Knowledge Tracing via Concept Aggregation and Relation-Based Attention

### v1

@@ V-D Visualization
5 Number Line, Mean, Probability of Two Distinct Events, Effect of Changing Dimensions of a Shape Prportionally, Estimation, Scale Factor, Greatest Common Factor, Recognize Linear Pattern
Visualization of student’s mastery. We also pivot towards gauging a student’s mastery of five concept clusters via visualization. The interaction sequence is derived from ASSIST09, consisting of the initial 50 interactions. A green rectangle implies that the student correctly answered the corresponding concept cluster while red rectangle means that the student failed. Student’s mastery of a concept cluster is calculated by his predicted performance on it. We find that our DGKT is capable of providing a coarse-grained but effective assessment despite the data scarcity issue.
Figure 7: A visualization of attention score from DGRKT and DGRKT -w/o RA reapectively.
Figure 8: The AUC result for diffarent $\lambda$ with different datasize.
Visualization of attention score. Here, we demonstrate the effectiveness of relation-based attention using a heatmap visualization. As shown in Fig. 7, we present the attention scores for four attention heads: the left side displays scores from DGRKT without relation-based attention (DGRKT -w/o RA), while the right side shows scores from DGRKT with relation-based attention. We observe that with the help of relation-based attention, DGRKT effectively distinguishes the relationships between different timesteps, focusing on those with closer connections. In contrast, the attention scores in DGRKT -w/o RA are smoother, indicating a reliance primarily on the sequential order.
Cluster analysis. We conducted an analysis on the aggregated concepts from ASSIST09, randomly selecting 8 concepts from each cluster. As shown in Table IV, it is evident that the first cluster primarily consists of fundamental and basic concepts, while the second cluster contains many advanced algebraic concepts. The third, fourth, and last clusters respectively include geometry concepts, advanced data representation, and statistical concepts. Although the concepts are not strictly categorized into five clusters, those with similar patterns in student interactions have been grouped together. These results clearly demonstrate the effectiveness of the concept aggregation process.
V-E Hyperparameter Analysis
@@
We conduct several experiment for $\lambda$ in concept representation of target domain. We present the auc of DGRKT with different $\lambda$ and different data size, shown in Fig. 8. When $\lambda$ is set too low, the constraints on the target domain become too strong, preventing the model from learning effective representations of the target domain, which leads to a decrease in target domain auc. Conversely, when $\lambda$ is set too high, the constraints diminish, causing the model to lose the information from its learned centroid embedding. As for $\lambda$, it achieves best performance when it’s set to 0.7.
VI Conclusion
@@
This paper introduces a novel approach, Domain Generalizable Knowledge Tracing (DGKT), as a solution to address the data scarcity issue within education systems. DGKT capitalizes on the utilization of multiple source domains to train a versatile KT network, enabling rapid adaptation to new target domains with commendable accuracy. Additionally, a novel concept aggregation technique along with target domain concept representation, enabling KT model to analyzes students’ interaction sequences from a coarse-grained perspective and effectively adapt to target domain with minimal data. Moreover, we propose domain-generalizable relation-based knowledge tracing (DGRKT) utilizing relation based attention encoder. Experimental evaluations conducted on five benchmark datasets demonstrate substantial improvements when compared to existing KT models. In conclusion, this study showcases the potential of DGKT in providing a versatile and accurate solution for knowledge tracing across a spectrum of educational domains.

### v2

@@ 4.4 Visualization
5 Number Line, Mean, Probability of Two Distinct Events, Effect of Changing Dimensions of a Shape Proportionally, Estimation, Scale Factor, Greatest Common Factor, Recognize Linear Pattern
4.5 Hyperparameter Analysis
@@
We conduct several experiments for parameter $\lambda$ (refereed to Eq. 7) in concept representation of target domain. We test the AUC of DGrKT with different values of $\lambda$ and different data sizes, shown in Fig. 7. When $\lambda$ is set too low, the constraints on the target domain become too strong, preventing the model from learning effective representations of the target domain, which leads to a decrease of AUC in target domain. Conversely, when $\lambda$ is set too high, the constraints diminish, causing the model to lose the information from its learned centroid embedding. As for $\lambda$, it achieves best performance when it is set to 0.7.
5 Conclusion and Discussion
@@
This paper introduces a novel approach, Domain Generalizable Knowledge Tracing framework, as a solution to address the data scarcity issue within education systems, which capitalizes on the utilization of multiple source domains to train a versatile KT network, enabling rapid adaptation to new target domains with commendable accuracy. Additionally, we propose the concept prototype to unify concepts from various domains via concept aggregation. Moreover, we propose a domain-generalizable relation-aware knowledge tracing (DGrKT) utilizing relation-aware attention encoder. Experimental evaluations conducted on five benchmark datasets demonstrate substantial improvements when compared to existing KT models. In conclusion, this study showcases the potential of DGKT in providing a versatile and accurate solution for knowledge tracing across a spectrum of educational domains.
Future research could explore the integration of additional semantic information. The concept aggregation process is purely data-driven, relying solely on student interaction data without integrating richer contextual information. For instance, incorporating textual representations of questions or domain-specific knowledge could potentially enhance generalization across diverse educational settings. Furthermore, beyond knowledge tracing, DGKT could be extended to other applications within educational systems, including personalized learning path recommendation and automatic difficulty adjustment. These extensions would further capitalize on DGKT’s domain-generalizable capabilities, broadening its impact on intelligent education systems.